OpenStudy (anonymous):

solution of this equation and how to do it? 11m-8-6m=22

OpenStudy (whpalmer4):

Add 8 to both sides giving 11m - 8 + 8 - 6m = 22 + 8 Simplify to get 5m = 30 Divide both sides by 5 to get m = ?

OpenStudy (whpalmer4):

As long as you do the same thing on both sides of the equal sign, the equation continues to be true. Just shuffle and combine until you get the answer, that's all there is to algebra :-)

OpenStudy (anonymous):

thank you!